OpenStudy (mckenzieandjesus):

Solve the system using elimination. 3x – y = 28 3x + y = 14

OpenStudy (mckenzieandjesus):

(8, –4) (–7, 7) (7, –7) (–4, 8)

Vocaloid (vocaloid):

hint: add the two equations together

Vocaloid (vocaloid):

start with the x's 3x + 3x = ?

OpenStudy (mckenzieandjesus):

3+3=6 28+14=42

Vocaloid (vocaloid):

almost 3 + 3 = 6 3x + 3x = ___ fill in the blank

OpenStudy (mckenzieandjesus):

6x?

Vocaloid (vocaloid):

good! now go to the y's -y + y = ?

OpenStudy (mckenzieandjesus):

0?

Vocaloid (vocaloid):

good! now we have 6x = 42 x = ___ fill in the blank

OpenStudy (mckenzieandjesus):

do i add those together?

Vocaloid (vocaloid):

nope, just find x

Vocaloid (vocaloid):

6x = 42 what does x = ?

OpenStudy (mckenzieandjesus):

7

Vocaloid (vocaloid):

good! x = 7, so we look for the answer choice with an x-coordinate of 7 the only possibility is (7,-7)

OpenStudy (mckenzieandjesus):

yeah

OpenStudy (mckenzieandjesus):

can u help me with -5x+y=-5 -4x+2y=2

Vocaloid (vocaloid):

hint: multiply the first equation by 2 first

OpenStudy (mckenzieandjesus):

ok 10

Vocaloid (vocaloid):

nope -5x*2 = ?

OpenStudy (mckenzieandjesus):

-10?

Vocaloid (vocaloid):

don't forget the x

Vocaloid (vocaloid):

ok, the way I wrote it was a bit vague, but I meant (-5)*(2)*x = ?

OpenStudy (mckenzieandjesus):

-10x

OpenStudy (mckenzieandjesus):

what do I do next?

Vocaloid (vocaloid):

2*y = ?

OpenStudy (mckenzieandjesus):

whats y?

OpenStudy (texaschic101):

-5x + y = -5 ----multiply by -2 -4x + 2y = 2 ------------ 10x - 2y = 10 (result of multiplying by -2) -4x + 2y = 2 ------------add 6x = 12 -- divide by 6 on both sides x = 2 now sub 2 in for x in one of the original equations and find y. questions ?

OpenStudy (mckenzieandjesus):

-6?

OpenStudy (mckenzieandjesus):

-4*2+2=-6

OpenStudy (texaschic101):

-5x + y = -5 -5(2) + y = -5 -10 + y = -5 y = -5 + 10 y = 5 check... -4x + 2y = 2 -4(2) + 2(5) = 2 -8 + 10 = 2 2 = 2 (correct) so x = 2 and y = 5

OpenStudy (mckenzieandjesus):

someone told me 2,5 wasnt the right answer..

OpenStudy (texaschic101):

plug it into the original equations to check... -4x + 2y = 2 -5x + y = -5 -4(2) + 2(5) = 2 -5(2) + 5 = -5 -8 + 10 = 2 -10 + 5 = -5 2 = 2 (correct) 5 = 5 (correct) (2,5) is the solution because it checks out in both equations as correct

OpenStudy (mckenzieandjesus):

Thank You